# Are “Solved Issues” in SWE-bench Really Solved Correctly? An Empirical Study

### v1

@@ 4.4. RQ4: The Correctness of Suspicious Patches
Results: After manual inspection, each sampled patch is classified as incorrect, correct, or with uncertain correctness, as summarized in Table  5.
@@ 5.1. Implications
Handling under-specified issue statements. The result of RQ4 shows that a large proportion of suspicious patches have uncertain correctness due to under-specified requirements in the issue statement. These patches potentially introduce unintended or undesired behaviors, compromising the robustness of the target project. This calls for better issue-solving tools that are capable of detecting and refining under-specifeid requirements collaboratively with users.
Building a new benchmark with well-specified statements. Although the issue statements in SWE-bench are considered to be of high quality by human annotators (swe, [n. d.]), vague issue statements with under-specified requirements still exist, as shown in RQ4. Such issue statements can misguide issue-solving tools in generating incorrect patches, constraining the reliability of the benchmark. Practitioners should consider building better benchmarks, where issues are well-specified and leave less ambiguity.
@@ 8. Data Availability
Our code and data are available: https://doi.org/10.5281/zenodo.15030298

### v2

@@ 4.4. RQ4: The Correctness of Suspicious Patches
Results: After manual inspection, each sampled patch is classified as incorrect, correct, or with uncertain correctness, as summarized in Table  8.
@@ 5.1. Implications
Handling under-specified issue statements. The result of RQ4 shows that a large proportion of suspicious patches have uncertain correctness due to under-specified requirements in the issue statement. These patches potentially introduce unintended or undesired behaviors, compromising the robustness of the target project. This calls for better issue-solving tools that are capable of detecting and refining under-specified requirements collaboratively with users.
Building a new benchmark with well-specified statements. Although the issue statements in SWE-bench are considered to be of high quality by human annotators (swe, 2024), vague issue statements with under-specified requirements still exist, as shown in RQ4. Such issue statements can misguide issue-solving tools in generating incorrect patches, constraining the reliability of the benchmark. Practitioners should consider building better benchmarks, where issues are well-specified and leave less ambiguity.
@@ 6. Related Work
Software engineering agents
LLM-based agents promise to help automate various software engineering tasks (Roychoudhury et al., 2025). Agents for various tasks have been presented, such as fault localization (Kang et al., 2024; Xu et al., 2025; Jiang et al., 2025), issue solving (Yang et al., 2024; Zhang et al., 2024; Xia et al., 2025), automated program repair (Bouzenia et al., 2025; Cheng et al., 2025), repository setup (Bouzenia and Pradel, 2025b; Eliseeva et al., 2025; Milliken et al., 2025), and generating issue-reproducing tests (Mündler et al., 2024; Nashid et al., 2025). Researchers have started to study the behavior such agents to better understand their strengths and weaknesses (Bouzenia and Pradel, 2025a). Our work complements such efforts and will help improve future software engineering agents by critically analyzing the patches produced by the agents.
@@ 8. Data Availability
Our code and data are available: https://github.com/ZJU-CTAG/PatchDiff
9. Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62202420), Zhejiang Provincial Natural Science Foundation of China (No. LZ25F020003), the European Research Council (ERC, grant agreements 851895 and 101155832), and the German Research Foundation within the DeMoCo project.